OpenStudy (kramse18):

Now shw that if Q(x)->B as x->infintity, then any solution y(x) of equation (9) satisfies y(x)-> B/a as x-> infinity. equation 9: dy/dx+ay=Q(x)

OpenStudy (irishboy123):

had a go and thought i'd post as no-one else has had a try, but the integration constant keeps getting in the way \(y' + a y = Q(x)\) Using integrating factor \(y = \dfrac{\int e^{ax} Q(x) dx + C}{e^{ax}}\) IBP \(\int e^{ax} Q(x) dx \) \(= \int \dfrac{d}{dx}(\dfrac{1}{a}e^{ax}) Q(x) dx\) \(= \dfrac{1}{a}e^{ax} Q(x) - \int \dfrac{1}{a}e^{ax} Q'(x) dx\) \(\lim\limits_{ x \to \infty} y = \lim\limits_{ x \to \infty} \dfrac{ \dfrac{1}{a}e^{ax} Q(x) - \int \dfrac{1}{a}e^{ax} Q'(x) dx +C}{e^{ax}}\) \(\lim\limits_{ x \to \infty} Q(x) = B \) \(=\dfrac{B}{a} - \lim\limits_{ x \to \infty} \dfrac{ \int \dfrac{1}{a}e^{ax} Q'(x) dx +C}{e^{ax}} \qquad \triangle\) \(\lim\limits_{ x \to \infty} Q(x) = B \implies \lim\limits_{ x \to \infty} Q'(x) = 0\) So \(\triangle\) becomes \(=\dfrac{B}{a} - \lim\limits_{ x \to \infty} \dfrac{ \int \dfrac{1}{a}e^{ax} \bullet 0 dx + C}{e^{ax}} \) \(=\dfrac{B}{a} - \lim\limits_{ x \to \infty} \dfrac{C}{e^{a x}}\) and that depends on the value of a, so no cigar looks too complicated an approach too

OpenStudy (irishboy123):

well you can create counter-examples that show that this only works if a > 0 so that Constant issue goes away but this is wrong: \( \lim\limits_{ x \to \infty} \dfrac{ \int \dfrac{1}{a}e^{ax} \color{red}{\bullet 0 }dx + C}{e^{ax}}\) it really depends on the form of function \(G'(x)\)

OpenStudy (kramse18):

Well thank you very much. This helps a lot